# Stability-Aware Training of Neural Network Interatomic Potentials with Differentiable Boltzmann Estimators

### v1

@@ A.10 Analysis of Energy and Forces Errors
Training runs with high learning rate and low QM loss coefficient achieve greater improvements in stable simulation time relative to a baseline model trained only on QM reference data (Figure 9a). However, these training runs also incur a greater increase in the Mean Absolute Error (MAE) of force prediction on a held out test dataset (Figure 9b). Due to training on a single structural observable, the observable-matching component of the StABlE objective is ill-posed: a NNIP which collapses simulations onto a sparse set of states exactly matching the reference observable would globally minimize the observable-matching loss and yield indefinitely stable simulations, while incurring a large QM/force error. As the learning rate of the StABlE Training procedure is increased, the optimization is increasingly pushed towards this degenerate mode. Increasing the weight of the QM objective counteracts this tendency. Consequently, a Pareto frontier arises between stability and force prediction accuracy (Figure 9c). Some settings of learning rate and QM loss coefficient are Pareto suboptimal, while choosing among the remaining combinations causes one to move along the frontier. Incorporating additional training observables, particularly those which are dynamical in nature (e.g velocity autocorrelation functions), could counteract the degeneracy and push the Pareto frontier outwards.
Finally, we note that the observed energy MAE increase on aspirin for our chosen combination of learning rate ($\alpha=0.001$) and loss coefficient ($\lambda=10$) is from 0.87 to 1.4 $\mathrm{kcal\,mol}^{-1}$, while DFT error for energies on MD17 can be as high as 2.3 $\mathrm{kcal\,mol}^{-1}$ [undefs]. Thus, some of the error in the NNIP predictions could be attributable to inaccuracies in the underlying DFT data.
A.11 Effect of Simulation Timestep on Stability
To demonstrate that the instability problem in NNIPs is not simply due to an inappropriate choice of simulation timestep, we perform simulations at various timesteps for each pair of system and model (Figure 10).
Figure 10: Effect of timestep on simulation stability. Reducing the timestep does not reliably lead to an improvement in stability, and can sometimes worsen stability.
We observe that stability does not necessarily improve as the timestep is reduced. In fact, for some systems, stability seems to consistently decrease as the timestep is reduced. Similar behavior has been observed in neural network based solvers for ordinary differential equations [undefak].
A.12 Velocity Autocorrelation Function of Aspirin

### v2

@@ A.11 Analysis of Energy and Forces Errors
Training runs with high learning rate and low QM loss coefficient achieve greater improvements in stable simulation time relative to a baseline model trained only on QM reference data (Supplementary Figure 10a). However, these training runs also incur a greater increase in the Mean Absolute Error (MAE) of force prediction on a held out test dataset (Supplementary Figure 10b). Due to training on a single structural observable, the observable-matching component of the StABlE objective is ill-posed: a MLFF which collapses simulations onto a sparse set of states exactly matching the reference observable would globally minimize the observable-matching loss and yield indefinitely stable simulations, while incurring a large QM/force error. As the learning rate of the StABlE Training procedure is increased, the optimization is increasingly pushed towards this degenerate mode. Increasing the weight of the QM objective counteracts this tendency. Consequently, a Pareto frontier arises between stability and force prediction accuracy (Supplementary Figure 10c). Some settings of learning rate and QM loss coefficient are Pareto suboptimal, while choosing among the remaining combinations causes one to move along the frontier. Incorporating additional training observables, particularly those which are dynamical in nature (e.g., velocity autocorrelation functions), could counteract the degeneracy and push the Pareto frontier outwards.
Finally, we note that the observed energy MAE increase on aspirin for our chosen combination of learning rate ($\alpha=0.001$) and loss coefficient ($\lambda=10$) is from 0.87 to 1.4 $\mathrm{kcal\,mol}^{-1}$, while DFT error for energies on MD17 can be as high as 2.3 $\mathrm{kcal\,mol}^{-1}$ Faber et al. (2017). Thus, some of the error in the MLFF predictions could be attributable to inaccuracies in the underlying DFT data.
A.12 Effect of Simulation Timestep on Stability
We perform 100 ps simulations using various timesteps for the aspirin and Ac-Ala3-NHMe tetrapeptide systems, using a SchNet and NequIP potential respectively (Supplementary Figure 11).
Figure 11: Effect of timestep on simulation stability. Reducing the timestep does not completely eliminate instability, and can sometimes worsen stability.
We observe that instability is not completely eliminated as the timestep is reduced. For the tetrapeptide system, stability consistently decreases as the timestep is reduced. Similar behavior has been observed in neural network based solvers for ordinary differential equations Krishnapriyan et al. (2023). For the aspirin system, stability improves as the timestep is reduced, but does so very slowly (simulations are not completely stable even with a timestep of 0.05 fs , which is 10 times lower than the original timestep).
A.13 Velocity Autocorrelation Function of Aspirin